OpenStudy (anonymous):

A crate with mass 32.5 kg initially at rest on a warehouse floor is acted on by a net horizontal force of 140 N. a)What acceleration is produced? b)How far does the crate travel in 10.0 s? C)What is its speed at the end of 10.0 s?

OpenStudy (anonymous):

1)f=ma 140=32.5a a=4.31m.s^-2 2)s= ut +1\2a(t^2) = 0 + 1\2(4.31)(100) = 215.5m 3) v^2 = u^2 + 2as v = [sqrt{2(4.31)(215.5}] =43.1m.s^-1